PartnerObserving Log for 2007 Mar 20/21

Observers: Rick Pogge, Eichii Egami
Support Scientist: none
TO: Aaron Ceranski

The Vernal Equinox is at 1707 MST today. Bands of thick cloud, with blue between, have been coming over all afternoon. Might get an extended break around sunset tonight, but the prospects for later tonight are not good.

All times below are UTC 2007 March 21

0130 - We're getting a break in the clouds, so it looks good for observing. The engineering crew had the counterweights off to work on the right-back hydrostatic bearing today and are still having problems. We could be down for a few hours until they finish, re-install the counterweights, and re-balance the telescope. I have not yet been able to take the afternoon calibrations as a result.

0200 - There is a problem with the wiring of one of the sensors on the hydrostatic bearing that is preventing them from finishing and getting us online. It could take an hour or so.

0330 - Back online, opening the dome and initializing the telescope. Wind is up, so keeping the vent doors closed to keep the oil spray to a minimum.

0350 - First target: Dwarf/LeoT - GUI locked up on importing the OB. None of the usual non-drastic tricks got it back, so we restarted the lbc service on lbccontrol, then restarted the camera system.

0402 - Let's try that again... Using Gfastextra to collimate/focus the primary mirror... After big defocus, had to put in +1000 Z11 by hand to see the central obscuration, and then another +500 of Z11, and -1500 Z7 before lpcfpia could take over... averaging about 2min/iteration (mostly spent waiting on readout and file transfer).

0434 - Dwarf/LeoT - running OB LeoT_g - no clouds except on horizon, (but non-photometric due to thin high stuff), Winds ~10m/s, RH=36%, T=2.6C
  • Image1: 043556 - 0.9" (+), very round but not perfect (4pix FWHM)
  • Image2: 044140 - 0.9" & fairly round
  • Image3: 044732 - 0.9" and round, satellite streak crossing chips 1-3
  • Image4: 045337 - 1.1" and a little elongated, seeing puffed up a bit

0459 - checking primary with Rfastextra, then OB LeoT_R
  • Image1: 050754 - guider failed almost immediately - images elongated exposure ruined
  • Image2: 051348 - guiding OK - 0.9" and fairly round
  • Image3: 051947 - guiding OK - 1" and round
  • Image4: 052551 - guiding OK - 1.1+ (5pix) and reasonably round - seeing is starting to get scruffy watching the guide star readout

0533 - Zaritsky/NGC3351 - OB a high priority target, checking seeing after focus check, offset +5arcmin to get away from galaxy. 5s shot to check seeing. Good (0.8) so proceeding, running OB NGC_3351_V
  • Image01: 055033 - 0.8" and very round. Nice image.
  • Image02: 055410 - 0.9" and very round
  • Image03: 055756 - 0.9" and very round
  • Image04: 060148 - 0.9" and very round - guider and telescope are behaving themselves...
  • Image05: 060545 - 0.8" and very round
  • Image06: 060928 - 0.85" and very round
  • Image07: 061257 - <0.8" and very round
  • Image08: 061642 - 0.72" and very round
  • Image09: 062023 - 0.8" and very round
  • Image10: 062422 - 0.8" and very round
Notes Transfer queue stalled for ~6min at 0614UTC, some images arrived out of order (e.g., 061642 before 061257). Wind 6-8m/s and occasionally gusty, but images look good overall. Minor GUI hangs at various mid-exposure times.

0628 - Doing Bfastextra to check, then running OB NGC_3351_U. Had to offset N 5arcmin. Clouds starting to appear in west, may start to encroach soon
  • Image01: 063746 - guider failed, ran open loop - images elongated, 1.4"?
  • Image02: 064134 - guider failed, ran open loop - 1.1" FWHM and a little elongated.
  • Image03: 064526 - gs in 16s loop, star looks good, lost lock about half-way through - 1.2" but not very round
  • Image04: 064914 - gs in 8s loop, guided OK - seeing starting to degrade (wind is picking up, too), 1.2" and more round
  • Image05: 065303 - gs in 8s loop, guided OK - 1.3" and fairly round
  • Image06: 065634 - gs in 8s loop, guided OK - 1.2" and round
  • Image07: 070002 - gs in 16s loop, guided OK - 1.1" and round
  • Image08: 070339 - gs in 8s loop, guided OK - 1.5" seeing - continuing to get worse
  • Image09: 070722 - guider failed after ramp to 32s, ran open loop - 1.6" FWHM
  • Image10: 071109 - gs in 8s loop, guided OK - 1.6" images - ick

0720 - Clouds starting to encroach on our location. Pretty opaque stuff. Wind now 10-13m/s, T=2.3C, RH=40%

0740 - Mostly overcast, parking scope and watching for clearing

0900 - Some improvement, though lots of junk around.

0910 - Wind gusts above limit, closing up and waiting it out. Crew outside working with the Lowell DIMM report seeing estimate >2", so even if we were open it would be bad seeing.

1030 - Still closed, winds not improving, satellite shows greater threat of clouds towards the west/southwest

1040 - Taking bias and dark calibrations missing from earlier...

1100 - Conditions will not allow us to acquire twilight flats, so we're calling it a night, and an observing run.

Night Summary

Late start due to work on one of the hydraulic pads. Clouds mostly cleared away at sunset except around the horizon and to the far west.

Observed Targets:
  • Dwarf/LeoT - taken in 0.9-1" seeing, observation set completed, but one image of the r set lost to guider failure. Tried a restart, same failure so aborted, last image showed seeing a little scruffier than at the start.

  • Zaritsky/NGC3351 - managed to get V image set in 0.8-0.9" seeing, but through the U set we had problems with first the autoguider, then deteriorating seeing.

Shutdown by clouds for a while, reopened briefly to try to setup on a field to assess the seeing but got shutdown by increasing winds when gusts began to go over the safe limits. Winds never diminished, we finally decided to call it quits.

-- RichardPogge - 20 Mar 2007
Topic revision: r6 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback